# Provably Extracting the Features from a General Superposition

### v1

@@ 3.2 Estimating the Fourier Mass
Again to address integrability issues from restricting strictly to a hyperplane, we will add some small thickness by convolving with a Gaussian. Formally, for any function $g$, vector $v\in{\mathbb{R}}^{d}$ and $A\succeq 0$, we define
$I^{*}_{g}(v,A)\mathrel{\mathop{:}}=\int_{{\mathbb{R}}^{d}}\left|\widehat{g}(y)\right|^{2}e^{-(y-v)^{\top}A(y-v)}\,dy\,.$ (1)
To simulate a hyperplane orthogonal to say $b_{1},\dots,b_{k}$, we set $A=C(b_{1}b_{1}^{\top}+\dots+b_{k}b_{k}^{\top})$ for some sufficiently large $C$ (for technical reasons later on, we will need different values of $C$ for different coordinates to bound the branching in the search algorithm).
Now our goal will be to estimate $I^{*}_{f^{(\ell)}}(v,A)$ for various $v,A$. A simple, but critical observation, is that by substituting the definition
@@
Now setting $g(x)=f^{(\ell)}(x)=f(x)e^{-\left\lVert x\right\rVert^{2}/(2\ell^{2})}$, we can write the above as an expectation over both $x$ and $\Delta$ involving $f(x)f(x+\Delta)$ — and we can estimate this by sampling and querying the values $f(x),f(x+\Delta)$ (see Algorithm 1).
@@ 3.3 Bounding the Search Algorithm
First we discuss bounding the total number of candidates that our algorithm recurses on. By choosing the orthonormal basis $b_{1},\ldots,b_{d}$ randomly, we can ensure with high probability, once we fix the first two coordinates, any hyperplane orthogonal to both $b_{1},b_{2}$ that is not too close to the origin intersects at most one of the lines $\{tv_{i}\}_{t\in{\mathbb{R}}}$.
We will simply brute-force over a sufficiently fine grid for the first two coordinates. Now consider a candidate for the first $k$ coordinates $\alpha_{1},\ldots,\alpha_{k}$ for $k\geq 2$. We then probe candidates for $\alpha_{k+1}$ by evaluating $I^{\star}_{f^{(\ell)}}(v,A)$ at
@@
with $K_{1}=K_{2}=K_{k+1}=C_{2},K_{3}=\cdots=K_{k}=C_{1}$ with $C_{2}\gg C_{1}$. To see why we need different scales for different coordinates, ignore the first two coordinates (since we are brute forcing over these). For the remaining coordinates, we search over a finer grid, of width $\approx\frac{1}{\sqrt{C_{2}}}$. If we included all candidates for which the mass $I^{\star}_{f^{(\ell)}}(v,A)$ is nontrivial, then this could blow up the number of candidates by a constant factor (since e.g. both of the grid points on either side of the hidden Fourier spike would be valid candidates). Instead we choose only a sufficiently separated set of candidates. Then once we fix a coordinate, we need to relax the width of the Gaussian in that direction to $\approx\frac{1}{\sqrt{C_{1}}}$ to ensure that we lose only a small fraction of the Fourier mass (see Algorithm 2 for details).
@@
Finally, we discuss how we guarantee that our algorithm recovers all relevant directions. Indeed, to recover a direction $v_{i}$, we need $\widehat{f}(tv_{i})$ to be nontrivial for some $t$ bounded away from 0. In fact, we need $t\gg 1/\ell$ since the “tubes” on which $\widehat{f}$ is concentrated now have width $\approx 1/\ell$.
@@
Ensuring this is not totally trivial. As an example, if $\sigma_{i}(\cdot)$ were a polynomial function like $x^{3}$, then actually this need not be true. For large $\ell$, we can easily compute that the Fourier transform of $x^{3}e^{-\left\lVert x\right\rVert^{2}/(2\ell^{2})}$ decays exponentially with width scale $1/\ell$.
@@
How can we rule out such examples? Because the functions $\sigma_{i}(\cdot)$ are Lipschitz, we can choose $\ell$ sufficiently large so that the only way $\sigma_{i}(\cdot)$ can be close to a polynomial is if it is linear. For $\sigma_{i}(\cdot)$ that are close to linear, there is an obvious lack of identifiability since we could easily have two different sets of vectors $\{v_{1},\dots,v_{s}\}$ and $\{v_{1}^{\prime},\dots,v_{s}^{\prime}\}$ such that $a_{1}v_{1}+\dots+a_{s}v_{s}=a_{1}v_{1}^{\prime}+\dots+a_{s}v_{s}^{\prime}$ which of course implies that for all $x$,
$a_{1}v_{1}^{\top}x+\dots+a_{s}v_{s}^{\top}x=a_{1}{v_{1}^{\prime}}^{\top}x+\dots+a_{s}{v_{s}^{\prime}}^{\top}x\,.$
@@
Fortunately, this does not affect our ability to recover the overall function since we can learn the “linear part” separately. For $\sigma_{i}(\cdot)$ that are not too close to linear, we actually show that there must be some $t\gtrsim 1/\sqrt{\ell}\gg 1/\ell$ such that $\widehat{\sigma_{i}}(t)$ is non-negligible (see Lemma 4.9). This can then be used to argue that our algorithm must recover the corresponding direction $v_{i}$.
Once we recover the directions $v_{i}$, we can recover the functions by interpolating in Fourier space. The main bound for this part is 4.7, which we use to bound the error from interpolating the Fourier transform only on a discretization of a bounded interval instead of over all of $\mathbb{R}$}.
@@ 3.4 Organization
In Section 4, we present some general bounds on functions and their Fourier transforms that will be used in the analysis. Then in Section 5, we present our machinery for estimating the Fourier mass $I^{*}_{f^{(\ell)}}(v,A)$. In Section 6, we then present our algorithm that makes use of this machinery to locate the hidden directions $v_{i}$. In Section 7, we then show how to recover the function $f$ given estimates for the directions $v_{i}$. Finally in Section 8, we put everything together to prove our main theorems.
@@ 5 Estimating Fourier Weight
$I^{*}_{g}(v,A)\mathrel{\mathop{:}}=\int_{{\mathbb{R}}^{d}}\left|\widehat{g}(y)\right|^{2}e^{-(y-v)^{\top}A(y-v)}\,dy,$
Crucial to estimating this quantity is the following formula that rewrites $I^{*}_{g}(v,A)$ as a quadratic form in $g$ itself, rather than its Fourier transform. This will allow us to estimate $I^{*}_{g}(v,A)$ via sampling. The crucial point about this formula is that it can be interpreted as an integral over $x$, of an expectation over a distribution of $\Delta$, of $g(x)g(x+\Delta)$ times a unit complex rotation, and the distribution is independent of $g$.
@@
$I^{*}_{g}(v,A)=\int_{{\mathbb{R}}^{d}}\left(e^{-\mathrm{i}v^{\top}\Delta}\int_{{\mathbb{R}}^{d}}g(x+\Delta)g(x)\,dx\right)\,N_{0,2A}(\Delta)\,d\Delta\,.$
@@
Recall that in our original learning setup, we have query access to a function $f_{\sim}:{\mathbb{R}}^{d}\to{\mathbb{R}}$ that is close to $f$ in $L_{\infty}$ norm. We now present our algorithm that makes use of 5.2 to estimate $I^{*}_{f^{(\ell)}}(v,A)$ using queries to $f_{\sim}$. In order to implement this, there is one additional wrinkle that we now explain. Since 5.2 requires integrating $g(x)g(x+\Delta)$ over all $x$, we will set $g$ to be $f^{(\ell)}$ to ensure integrability. Because we have query access to $f_{\sim}$, which is close to $f$, we can then rewrite the integral over all $x$ of $f^{(\ell)}(x)f^{(\ell)}(x+\Delta)$ as an expectation over $x$ drawn from an appropriate Gaussian distribution of $f(x)f(x+\Delta)$, which we can then estimate by sampling.
@@ 11      Set $c_{j}\leftarrow\phi\cdot u\cdot w$
$\left|\,I-I^{*}_{f^{(\ell)}}(v,A)\right|\;\leq\;8\,(\pi\ell^{2})^{\frac{d}{2}}\!\left(\varepsilon+\sqrt{\frac{2\log(8/\delta)}{m}}\right).$
@@
$\begin{split}C\operatorname{\mathbb{E}}[c_{j}^{*}]&=\int\int f\left(Z-\frac{1}{2}\Delta\right)f\left(Z+\frac{1}{2}\Delta\right)e^{-\tfrac{\|\Delta\|^{2}}{4\ell^{2}}-\mathrm{i}v^{\top}\Delta}(\pi\ell^{2})^{\frac{d}{2}}N_{0,\frac{\ell^{2}}{2}I_{d}}(Z)\cdot N_{0,2A}(\Delta)dZd\Delta\\ &=\int\int f^{(\ell)}\left(Z-\frac{1}{2}\Delta\right)f^{(\ell)}\left(Z+\frac{1}{2}\Delta\right)e^{-\mathrm{i}v^{\top}\Delta}N_{0,2A}(\Delta)dZd\Delta\\ &=I^{*}_{f^{(\ell)}}(v,A).\end{split}$
@@
$\Big|\operatorname{\mathbb{E}}[I]-I^{*}_{f^{(\ell)}}(v,A)\Big|\leq C\,\varepsilon(2+\varepsilon).$

### v2

@@ 3.2 Estimating the Fourier Mass
Again, to ensure that the notion of “total weight on a hyperplane” is well-defined, we need to address integrability. Rather than restricting strictly to a hyperplane, we will add some small Gaussian thickening. Formally, for any function $g$, vector $v\in{\mathbb{R}}^{d}$ and $A\succeq 0$, we define
$I^{\star}_{g}(v,A)\mathrel{\mathop{:}}=\int_{{\mathbb{R}}^{d}}\left|\widehat{g}(y)\right|^{2}e^{-(y-v)^{\top}A(y-v)}\,dy\,.$ (1)
To simulate a hyperplane orthogonal to say $b_{1},\dots,b_{k}$, we set $A=C(b_{1}b_{1}^{\top}+\dots+b_{k}b_{k}^{\top})$ for some sufficiently large $C$ (for technical reasons later on, we will need different values of $C$ for different coordinates to bound the branching in the search algorithm). This choice of $A$ ensures that the integral is dominated by the Fourier mass on points that are close to the hyperplane through $v$, orthogonal to $b_{1},\dots,b_{k}$.
Now our goal will be to estimate $I^{\star}_{f^{(\ell)}}(v,A)$ for various $v,A$. A simple, but critical observation, is that by substituting the definition
@@
Now setting $g(x)=f^{(\ell)}(x)=f(x)e^{-\left\lVert x\right\rVert^{2}/(2\ell^{2})}$, we can write the above as an expectation over both $x$ and $\Delta$ involving values of $f$ itself:
$I^{\star}_{f^{(\ell)}}(v,A)=(\pi\ell^{2})^{\frac{d}{2}}\operatorname{\mathbb{E}}_{x\sim\mathcal{N}(0,\frac{\ell^{2}}{2}I_{d}),\Delta\sim\mathcal{N}(0,2A)}\!\Big[e^{-\frac{\|\Delta\|^{2}}{4\ell^{2}}-iv^{\top}\Delta}\cdot f\!\left(x-\tfrac{1}{2}\Delta\right)\,f\!\left(x+\tfrac{1}{2}\Delta\right)\Big]\,.$ (2)
Since $f(x)$ is bounded, the quantity inside the expectation is bounded, and we can estimate the expectation by sampling and querying the values $f\left(x-\tfrac{1}{2}\Delta\right),f\left(x+\tfrac{1}{2}\Delta\right)$ for many $x,\Delta$ (see Algorithm 1). This is the key subroutine that allows us to estimate the Fourier mass on various hyperplane-like slices. Stated more concisely:
Corollary 3.1 (Informal, see Corollary 5.3).
We can estimate $I^{\star}_{f^{(\ell)}}(v,A)$ to within $\varepsilon\cdot(\pi\ell^{2})^{\frac{d}{2}}$ additive error with high probability using polynomially many queries.
@@
Note that the scaling $(\pi\ell^{2})^{\frac{d}{2}}$ is the “right” scaling. To see this intuitively, if $|f(x)|=1$ everywhere, then $\int_{{\mathbb{R}}^{d}}|\widehat{f^{(\ell)}}(y)|^{2}dy=\int_{{\mathbb{R}}^{d}}|f^{(\ell)}(x)|^{2}dx=(\pi\ell^{2})^{\frac{d}{2}}$. Estimating the inner expectation in (2) to within $\varepsilon$ additive error thus roughly corresponds to an additive error of $\varepsilon$-fraction of the total Fourier mass. We will always aim to preserve at least an inverse polynomial fraction of the Fourier mass when we localize, so we will only ever need to estimate $I^{\star}_{f^{(\ell)}}(v,A)$ to within inverse polynomial (times $(\pi\ell^{2})^{\frac{d}{2}}$) accuracy.
@@ 3.3 Bounding the Search Algorithm
Recall the Fourier mass of $f^{(\ell)}$ is concentrated on tubes centered around the lines $\{tv_{i}\}_{t\in{\mathbb{R}},i\in[n]}$. First we ensure that the projections of the lines $\{tv_{i}\}_{t\in{\mathbb{R}},i\in[n]}$ onto the plane formed by the first two coordinates $b_{1},b_{2}$ are still separated lines. To achieve this, we can simply choose the orthonormal basis $b_{1},\ldots,b_{d}$ randomly. A direct anti-concentration argument and union bound shows that then, the projected lines are separated with high probability (see 6.5 for details).
This means that once we fix the first two coordinates $(\alpha_{1},\alpha_{2})$, as long as $(\alpha_{1},\alpha_{2})$ is bounded away from the origin, at most one of the lines $tv_{i}$ can still be close to any point whose first two coordinates are $(\alpha_{1},\alpha_{2})$.
Bounding the Number of Candidates
We now discuss bounding the total number of candidates that our algorithm recurses on. We will simply brute-force over a sufficiently fine grid for the first two coordinates (with a neighborhood around the origin cut out). The number of candidates for the first two coordinates is then polynomially bounded. We will focus on ensuring that the number of candidates does not blow up as we search over the remaining coordinates.
Consider a candidate for the first $k$ coordinates $\alpha_{1},\ldots,\alpha_{k}$ for $k\geq 2$. We then probe candidates for $\alpha_{k+1}$ by evaluating $I^{\star}_{f^{(\ell)}}(v,A)$ at
@@
with $K_{1}=K_{2}=K_{k+1}=C_{2},K_{3}=\cdots=K_{k}=C_{1}$ with $C_{2}\gg C_{1}$. To see why we need different scales for different coordinates, ignore the first two coordinates (since we are brute forcing over these). For the remaining coordinates, when searching over say the k+1k+1st coordinate $\alpha_{k+1}$, we search over a finer grid, of width $\approx\frac{1}{\sqrt{C_{2}}}$.
Initially, we view each value on this grid as a candidate but then we filter down to only the candidates for which the mass $I^{\star}_{f^{(\ell)}}(v,A)$ is nontrivial where $v=\sum_{j\leq k+1}\alpha_{j}b_{j}$. At this point, if we kept and recursed on all of the remaining candidates, then this could blow up the number of candidates by a constant factor (since e.g. both of the grid points on either side of the hidden Fourier spike would be valid candidates). Instead we choose only a sufficiently separated set of candidates. Naively, this runs into the potential issue that we could be losing a constant fraction of the Fourier mass if the true spike is between two candidate grid points.
The fix is that after probing with a fine grid and then pruning, we relax the width of the Gaussian for all steps afterward. Once we choose a value for the k+1k+1st coordinate, we can relax the width of the Gaussian in that direction to $\approx\frac{1}{\sqrt{C_{1}}}$ to retain all but a negligible fraction of the Fourier mass from the previous iteration of the recursion. Stated more precisely, for
$A_{k}=\sum_{j\leq k}K_{j}\,b_{j}b_{j}^{\top},A_{k+1}=\sum_{j\leq k+1}K_{j}\,b_{j}b_{j}^{\top}$
where $K_{1}=K_{2}=C_{2},K_{3}=\cdots=K_{k}=K_{k+1}=C_{1}$, we ensure that $I^{\star}_{f^{(\ell)}}(v,A_{k})-I^{\star}_{f^{(\ell)}}(v,A_{k+1})$ is sufficiently small. See Algorithm 2 for the full details of the search algorithm.
Recall that after fixing the first two coordinates $(\alpha_{1},\alpha_{2})$, there is at most one line $tv_{i}$ for $i\in[n]$ that can be close. This essentially uniquely determines the remaining coordinates. Putting everything together, we have:
Claim 3.3 (Informal, see 6.8).
Assuming that the projections of the lines $\{tv_{i}\}_{t\in{\mathbb{R}},i\in[n]}$ onto the plane formed by the first two coordinates $b_{1},b_{2}$ are separated, then for each choice for the first two coordinates away from the origin, the algorithm recurses on at most one candidate for each of the remaining coordinates. Thus, the total number of candidates that the algorithm recurses on is polynomially bounded.
Ensuring Completeness
Finally, we discuss how we guarantee that when we run the above search algorithm, we recover all relevant directions. For each branch, we recurse until we have chosen all $d$ coordinates $(\alpha_{1},\dots,\alpha_{d})$ (or the branch terminates with no valid candidates). The collection of points that we recover at the end gives us a collection of directions. Just from the fact that the Fourier transform $\widehat{f^{(\ell)}}$ is concentrated on tubes around the lines $\{tv_{i}\}_{t\in{\mathbb{R}},i\in[n]}$, we can argue that the recovered directions must all be close to some true direction $v_{i}$.
@@
We now discuss how to ensure that this collection is complete i.e. recovers all of the directions $v_{i}$. To recover a direction $v_{i}$, it suffices to show that there is a non-negligible amount of Fourier mass associated with the line $\{tv_{i}\}$ at some scale bounded away from 0. In fact, we need $t\gg 1/\ell$ since the “tubes” around the lines on which $\widehat{f^{(\ell)}}$ is concentrated now have width $\approx 1/\ell$. The main algorithmic statement is:
Claim 3.4 (Informal, see 6.7).
If for some direction $v_{i}$, there exists $t\gg 1/\ell$ such that the Fourier mass of $f^{(\ell)}$ around $tv_{i}$ is at least an inverse polynomial fraction of the total Fourier mass, then the search algorithm will recover a direction close to $v_{i}$.
Intuitively, this claim holds because if the Fourier mass around some point $tv_{i}$ is nontrivial for some $t\gg 1/\ell$, then one of the candidates that we search over will be close to $tv_{i}$ and this will ensure that we recover the direction $v_{i}$. We now proceed with this intuitive understanding (see the proof of 6.7 for a more formal treatment).
@@
The previous condition ends up being essentially equivalent to finding some $t\gg 1/\ell$ such that $\widehat{\sigma_{i}^{(\ell)}}(t)$ is nontrivial. While the needed condition seems intuitive for natural function classes, ensuring it requires additional work. As an example, if $\sigma_{i}(\cdot)$ were a polynomial function like $x^{3}$, then actually this need not be true. For large $\ell$, we can easily compute that the Fourier transform of $\sigma_{i}^{(\ell)}(x)=x^{3}e^{-\left\lVert x\right\rVert^{2}/(2\ell^{2})}$ decays exponentially with width scale $1/\ell$.
@@
How can we rule out such examples? Because the functions $\sigma_{i}(\cdot)$ are Lipschitz, we can choose $\ell$ sufficiently large so that the only way $\sigma_{i}(\cdot)$ can be close to a polynomial is if it is linear. For $\sigma_{i}(\cdot)$ that are close to linear, there is an obvious lack of identifiability since we could easily have two different sets of vectors $\{v_{1},\dots,v_{s}\}$ and $\{v_{1}^{\prime},\dots,v_{s}^{\prime}\}$ such that $v_{1}+\dots+v_{s}=v_{1}^{\prime}+\dots+v_{s}^{\prime}$ which of course implies that for all $x$,
$v_{1}^{\top}x+\dots+v_{s}^{\top}x={v_{1}^{\prime}}^{\top}x+\dots+{v_{s}^{\prime}}^{\top}x\,.$
@@
Fortunately, this does not affect our ability to recover the overall function since we can learn the “linear part” separately. Crucially, we actually show that, in some sense, linear response functions are the only obstruction. For $\sigma_{i}(\cdot)$ that are not too close to linear, we show, via careful analysis, that there must be some $t\gtrsim 1/\sqrt{\ell}\gg 1/\ell$ such that $\widehat{\sigma_{i}^{(\ell)}}(t)$ is non-negligible (see Lemma 4.9). This can then be used to argue that our algorithm must recover the corresponding direction $v_{i}$. The identifiability of the directions $v_{i}$ for which the response function $\sigma_{i}(\cdot)$ is non-linear then follows as a consequence of this characterization.
3.4 Function Recovery
Once we recover the directions $v_{i}$, we can recover the function by interpolating in Fourier space. We estimate the Fourier transform $\widehat{f^{(\ell)}}(y)$ for $y$ on a bounded discrete grid on each of the lines $\{tv_{i}\}_{t\in{\mathbb{R}},i\in[n]}$. Then for each line, we apply the Fourier inversion formula, interpolating between the grid points and zeroing out the part outside the bounded grid. We show that this approximately recovers each of the ridge functions $\sigma_{i}(v_{i}^{\top}x)$, and thus adding up the reconstructions along each of the lines recovers $f$.
The main bound that we need is 4.7 which shows that for any bounded Lipschitz function over $\mathbb{R}$}, we can reconstruct it to within a small error by interpolating its Fourier transform on a bounded discrete grid. We then apply this bound in Lemma 7.4 to control the overall error in recovering the function $f$.
@@ 3.5 Organization
In Section 4, we present some general bounds on functions and their Fourier transforms that will be used in the analysis. Then in Section 5, we present our machinery for estimating the Fourier mass $I^{\star}_{f^{(\ell)}}(v,A)$. In Section 6, we then present our algorithm that makes use of this machinery to locate the hidden directions $v_{i}$. In Section 7, we then show how to recover the function $f$ given estimates for the directions $v_{i}$. Finally in Section 8, we put everything together to prove our main theorems.
@@ 5 Estimating Fourier Weight
$I^{\star}_{g}(v,A)\mathrel{\mathop{:}}=\int_{{\mathbb{R}}^{d}}\left|\widehat{g}(y)\right|^{2}e^{-(y-v)^{\top}A(y-v)}\,dy,$
Crucial to estimating this quantity is the following formula that rewrites $I^{\star}_{g}(v,A)$ as a quadratic form in $g$ itself, rather than its Fourier transform. This will allow us to estimate $I^{\star}_{g}(v,A)$ via sampling. The crucial point about this formula is that it can be interpreted as an integral over $x$, of an expectation over a distribution of $\Delta$, of $g(x)g(x+\Delta)$ times a unit complex rotation, and the distribution is independent of $g$.
@@
$I^{\star}_{g}(v,A)=\int_{{\mathbb{R}}^{d}}\left(e^{-\mathrm{i}v^{\top}\Delta}\int_{{\mathbb{R}}^{d}}g(x+\Delta)g(x)\,dx\right)\,N_{0,2A}(\Delta)\,d\Delta\,.$
@@
Recall that in our original learning setup, we have query access to a function $f_{\sim}:{\mathbb{R}}^{d}\to{\mathbb{R}}$ that is close to $f$ in $L_{\infty}$ norm. We now present our algorithm that makes use of 5.2 to estimate $I^{\star}_{f^{(\ell)}}(v,A)$ using queries to $f_{\sim}$. In order to implement this, there is one additional wrinkle that we now explain. Since 5.2 requires integrating $g(x)g(x+\Delta)$ over all $x$, we will set $g$ to be $f^{(\ell)}$ to ensure integrability. Because we have query access to $f_{\sim}$, which is close to $f$, we can then rewrite the integral over all $x$ of $f^{(\ell)}(x)f^{(\ell)}(x+\Delta)$ as an expectation over $x$ drawn from an appropriate Gaussian distribution of $f(x)f(x+\Delta)$, which we can then estimate by sampling.
@@ 11      Set $c_{j}\leftarrow\phi\cdot u\cdot w$
$\left|\,I-I^{\star}_{f^{(\ell)}}(v,A)\right|\;\leq\;8\,(\pi\ell^{2})^{\frac{d}{2}}\!\left(\varepsilon+\sqrt{\frac{2\log(8/\delta)}{m}}\right).$
@@
$\begin{split}C\operatorname{\mathbb{E}}[c_{j}^{*}]&=\int\int f\left(Z-\frac{1}{2}\Delta\right)f\left(Z+\frac{1}{2}\Delta\right)e^{-\tfrac{\|\Delta\|^{2}}{4\ell^{2}}-\mathrm{i}v^{\top}\Delta}(\pi\ell^{2})^{\frac{d}{2}}N_{0,\frac{\ell^{2}}{2}I_{d}}(Z)\cdot N_{0,2A}(\Delta)dZd\Delta\\ &=\int\int f^{(\ell)}\left(Z-\frac{1}{2}\Delta\right)f^{(\ell)}\left(Z+\frac{1}{2}\Delta\right)e^{-\mathrm{i}v^{\top}\Delta}N_{0,2A}(\Delta)dZd\Delta\\ &=I^{\star}_{f^{(\ell)}}(v,A).\end{split}$
@@
$\Big|\operatorname{\mathbb{E}}[I]-I^{\star}_{f^{(\ell)}}(v,A)\Big|\leq C\,\varepsilon(2+\varepsilon).$